# Learning from Mistakes: Understanding Ad-hoc Logs through Analyzing Accidental Commits

### v1

@@ Abstract
Developers often insert temporary “print” or “log” instructions into their code to help them better understand runtime behavior, usually when the code is not behaving as they expected. Despite the fact that such monitoring instructions, or “ad-hoc logs,” are so commonly used by developers, there is almost no existing literature that studies developers’ practices in how they use them. This paucity of knowledge of the use of these ephemeral logs may be largely due to the fact that they typically only exist in the developers’ local environments and are removed before they commit their code to their revision control system. In this work, we overcome this challenge by observing that developers occasionally mistakenly forget to remove such instructions before committing, and then they remove them shortly later. Additionally, we further study such developer logging practices by watching and analyzing live-streamed coding videos. Through these empirical approaches, we study where, how, and why developers use ad-hoc logs to better understand their code and its execution. We collect 27 GB of accidental commits that removed 548,880 ad-hoc logs in JavaScript from GitHub Archive repositories to provide the first large-scale dataset and empirical studies on ad-hoc logging practices. Our results reveal several illuminating findings, including a particular propensity for developers to use ad-hoc logs in asynchronous and callback functions. Our findings provide both empirical evidence and a valuable dataset for researchers and tool developers seeking to enhance ad-hoc logging practices, and potentially deepen our understanding of developers’ practices towards understanding of software’s runtime behaviors.
@@
Empirical Software Engineering, Logs, Mining Software Repository
@@ I Introduction
In contrast, to the best of our knowledge, the second type of logs—ad-hoc logs —remains largely understudied in software-engineering literature [1]. Despite their essential role in software development, particularly in debugging [13, 14, 3, 15, 16, 17], the transient and local nature of ad-hoc logs has limited the research community’s ability to study and understand them.
Fortunately, for the purposes of research, not all ad-hoc logs are completely hidden in developers’ local environments. Like anyone, developers occasionally make mistakes when pushing code to version-control systems, ranging from simple, easily fixable errors [18] to unintentionally exposing sensitive information, such as API keys [19]. Our research leverages this serendipitous opportunity to address a significant research gap:
Approximate the comprehension of dynamic aspect of software by analyzing large-scaled empirical evidences of ad-hoc logs.
@@
For deepening our understanding of the use of these logs, we also qualitatively analyze the 36 hours of developers’ live-stream videos collected by Alaboudi et al. [13]. This further analysis allows us to see all of the temporary placement and use of their ad-hoc logs, instead of just the ones that were mistakenly committed to the version-control system. It also allows us to better understand the situations and comprehension tasks in which the developer needed to use them. Whereas the GitHub data gives us breadth and scalability, the live-stream data gives us depth and qualitative insight.
The results of our studies reveal a number of interesting findings. For example, we found a common logging behavior in which developers made efforts in structuring the console outputs so that they could pinpoint and compare the program states or executions on the fly.
@@ II Background
The challenges of working with human subjects for incremental studies led us to draw on previous mining software repositories (MSR) research, especially for those works that uncover hidden patterns in revision histories. For example, Beyer et al. identified bug-fixing patterns from GitHub histories [40], and Sinha et al. published data on secret key leaks caused by developers’ mismanagement of public version control systems [41]. Observing developers’ frequent mistakes in version control, we mined and published ad-hoc logs by examining patches made after inadvertent changes.
@@ IV Dataset Curation
Live-Streamed Videos. In this subsection, we detail our data-curation process, building on the work of Alaboudi et al. [14], where researchers observed developers live streaming their software development and documented their working cycles. Developers, in particular, are adopting live streaming to increase visibility, facilitate education, and foster community [43]. Such a dataset provides unobtrusive observation of behaviors in natural settings, which can reduce observer bias [44]. Despite its potential, live streaming remains underexplored in software-engineering research. Fortunately, prior work by Alaboudi et al. [14] has made significant strides by collecting and labeling live-streamed videos and developing the observer-dev.online platform [45], tailored for software-engineering researchers. We used all 15 videos (total duration: 36:01:30) that they curated and conducted a thematic analysis, with both the dataset and our qualitative findings made publicly available [46]. While little research has explored why and how developers insert log statements, Alaboudi et al. [14] segmented each video into debugging episodes, such as “Testing the program and reading the outputs” or “Interacting with a file of code (Edit).” These segments allowed us to efficiently pinpoint moments when developers began adding log statements to their source code.
@@ VII Discussion
While these findings may seem intuitive, the collected glints shed light on re-examining previous metrics and refining our understanding of software complexity. There has long been debate over which metrics—Cyclomatic Complexity, Cognitive Complexity, or even custom models—can best capture the difficulties perceived by developers [60]. However, as discussed before, human studies are costly. Future studies could leverage our empirical dataset to revisit these debates or to evaluate new complexity metrics. Similarly, computer-science educators and software maintainers can observe students’ or developers’ logging behaviors to approximate program complexity, helping them decide when to initiate refactoring or when to adjust the difficulty level.
@@
Not only do logs need to be placed thoughtfully and contain relevant content; their presentation must also sharpen each glimpse into the code’s layers. Finding 2 shows that many ad-hoc logs include labels and formatting strings, with developers often investing time to create custom debug functions for added clarity. Finding 3 further reveals that to understand or compare states and execution flows, developers carefully craft anchor strings and structure complex objects. Given the rapid pace of program execution relative to human perception, large volumes of log outputs often flood the screen quickly. Consequently, developers must map console outputs back to the source code (spatial mapping) and, in the context of asynchronous functions or third-party libraries, use anchors to mark specific execution timings (temporal mapping). While Log-it has made initial progress in supporting easier mapping [3], future tools could further enhance these mappings by incorporating prior works in visualizations [54] or recommendation systems [32].
@@
External Threats to Validity. The findings may not generalize to other languages, such as Java and Python. However, we provide scripts to facilitate similar research in different programming contexts. The external validity of the live-stream videos, including representativeness and selection criteria, has been discussed and addressed in the original research by Alaboudi et al. [13].
@@ IX Acknowledgment
We extend our gratitude to the anonymous reviewers for their thorough and insightful feedback. This research is supported by the Dieter Schwarz Stiftung Foundation.

### v2

@@ Abstract
Developers often insert temporary “print” or “log” instructions into their code to help them better understand runtime behavior, usually when the code is not behaving as they expected. Despite the fact that such monitoring instructions, or “ad-hoc logs,” are so commonly used by developers, there is almost no existing literature that studies developers’ practices in how they use them. This paucity of knowledge of the use of these ephemeral logs may be largely due to the fact that they typically only exist in the developers’ local environments and are removed before they commit their code to their revision control system. In this work, we overcame this challenge by observing that developers occasionally mistakenly forget to remove such instructions before committing, and then they remove them shortly later. Additionally, we further studied such developer logging practices by watching and analyzing live-streamed coding videos. Through these empirical approaches, we presented where, how, and why developers use ad-hoc logs to better understand their code and its execution. We collected 27 GB of accidental commits that removed 548,880 ad-hoc logs in JavaScript from GitHub Archive repositories to provide the first large-scale dataset and empirical studies on ad-hoc logging practices. Our results revealed several illuminating findings, including a particular propensity for developers to use ad-hoc logs in asynchronous and callback functions. Our findings provided both empirical evidence and a valuable dataset for researchers and tool developers seeking to enhance ad-hoc logging practices, and potentially deepen our understanding of developers’ practices towards understanding of software’s runtime behaviors.
@@
Empirical Software Engineering, Ad-hoc Logs, Mining Software Repository
@@ I Introduction
In contrast, to the best of our knowledge, the second type of logs—ad-hoc logs [^1] In this work, we use “log” and “logging statement” interchangeably to refer to a logging instruction, as in prior work [3].—remains largely understudied in software-engineering literature [1]. Despite their essential role in software development, particularly in debugging [13, 14, 3, 15, 16, 17], the transient and local nature of ad-hoc logs has limited the research community’s ability to study and understand them.
Fortunately, for the purposes of research, not all ad-hoc logs are completely hidden in developers’ local environments. Like anyone, developers occasionally make mistakes when pushing code to version-control systems, such as GitHub, ranging from simple, easily fixable errors [18] to exposing sensitive information, such as API keys [19]. Our research leverages this serendipitous opportunity to address a significant research gap: inferring developer challenges in understanding runtime behavior of code by analyzing large-scale empirical evidence of ad-hoc logs.
@@
For deepening our understanding of the use of these logs, we also qualitatively analyzed the 36 hours of developers’ live-stream videos collected by Alaboudi et al. [13]. This further analysis allows us to see all of the temporary placement and use of their ad-hoc logs, instead of just the ones that were mistakenly committed to the version-control system. It also allows us to better understand the situations and comprehension tasks in which the developer needed to use them. Whereas the GitHub data gives us breadth and scalability, the live-stream data gives us depth and qualitative insight.
The results of our studies revealed a number of interesting findings. For example, we found a common logging behavior in which developers made efforts in structuring the console outputs so that they could pinpoint and compare the program states or executions on the fly.
@@ II Background
The challenges of working with human subjects for incremental studies led us to draw on previous mining software repositories (MSR) research, especially for those works that uncover hidden patterns in revision histories. For example, Beyer et al. identified bug-fixing patterns from GitHub histories [40], and Sinha et al. published data on secret key leaks caused by developers’ mismanagement of public version control systems [41]. Observing developers’ frequent mistakes in version control, we mined and published ad-hoc logs by examining patches made after inadvertent changes. Although it is possible that developers could intentionally commit such logs, we consider them as mistakes for multiple reasons: (1) we cannot know the initial developer intent and the subsequent deletion may indicate a fix, (2) console.log instructions in JavaScript output to transient consoles for users, and (3) community guidelines (e.g., Airbnb [42]) and widely used linters (e.g., ESLint [43]) discourage committing console.log statements to version-control systems.
@@ IV Dataset Curation
Live-Streamed Videos. In this subsection, we detail our data-curation process, building on the work of Alaboudi and LaToza [14], where researchers observed developers live streaming their software development and documented their working cycles. Developers, in particular, are adopting live streaming to increase visibility, facilitate education, and foster community [45]. Such a dataset provides unobtrusive observation of behaviors in natural settings, which can reduce observer bias [46]. Despite its potential, live streaming remains underexplored in software-engineering research. Fortunately, prior work by Alaboudi and LaToza[14] has made significant strides by collecting and labeling live-streamed videos and developing the observer-dev.online platform [47], tailored for software-engineering researchers. We used all 15 videos (total duration: 36:01:30) that they curated and conducted a thematic analysis, with both the dataset and our qualitative findings made publicly available [48]. While little research has explored why and how developers insert log statements, Alaboudi and LaToza [14] segmented each video into debugging episodes, such as “Testing the program and reading the outputs” or “Interacting with a file of code (Edit).” These segments allowed us to efficiently pinpoint moments when developers began adding log statements to their source code.
@@ VII Discussion
While these findings may seem intuitive, the collected glints shed light on re-examining previous metrics and refining our understanding of software complexity. There has long been debate over which metrics—Cyclomatic Complexity, Cognitive Complexity, or even custom models—can best capture the difficulties perceived by developers [61, 62]. However, as discussed before, human studies are costly. Future studies could leverage our empirical dataset to revisit these debates or to evaluate new complexity metrics. Similarly, computer-science educators and software maintainers can observe students’ or developers’ logging behaviors to approximate program complexity, helping them decide when to initiate refactoring or when to adjust the difficulty level.
@@
Not only do logging statements need to be placed thoughtfully and contain relevant content; their presentation must also sharpen each glimpse into the call stacks or data structure. Finding 2 shows that many ad-hoc logs include labels and formatting strings, with developers often investing time to create custom debug functions for added clarity. Finding 3 further reveals that to understand or compare states and execution flows, developers carefully craft anchor strings and structure complex objects. Given the rapid pace of program execution relative to human perception, large volumes of log outputs often flood the screen quickly. Consequently, developers must map console outputs back to the source code (spatial mapping) and, in the context of asynchronous functions or third-party libraries, use anchors to mark specific execution timings (temporal mapping). While Log-it has made initial progress in supporting easier mapping [3], future tools could further enhance these mappings by incorporating prior works in visualizations [56] or recommendation systems [32].
@@
External Threats to Validity. The findings may not generalize to other languages, such as Java and Python. However, we provide scripts to facilitate similar research in different programming contexts. The external validity of the live-stream videos, including representativeness and selection criteria, has been discussed and addressed in the original research by Alaboudi and LaToza [13].
@@ IX Acknowledgment
We extend our gratitude to the anonymous reviewers for their thorough and insightful feedback. We would also like to thank Heng Du, Junling Wang, Xiaotian Su, Zeyu Xiong, Gustavo Umbelino, Behnood Masoudi, Monil Narang for providing feedback and support. This research is partially supported by the Dieter Schwarz Stiftung Foundation.